OpenStudy (richyw):

how would I show that this indefinite integral equals 1/2?

OpenStudy (richyw):

\[\lim_{r\rightarrow \infty}\int^r_0p^3e^{-p^2}=\frac{1}{2} \]

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

integration by parts will work here...i think...lets work on that

OpenStudy (richyw):

tried that

OpenStudy (richyw):

after three times, to get eliminate p, it just seems to get uglier

OpenStudy (turingtest):

\[u=e^{-p^2}\implies du=-2pe^{-p^2}dp\]\[\ln u=-p^2\]\[p^3e^{-p^2}dp=-\frac12p^2du=\frac12\ln udu\]maybe?

OpenStudy (turingtest):

just trying things...

OpenStudy (turingtest):

now you could integrate by parts I think

OpenStudy (anonymous):

hmm... \(\frac{1}{2}(u ln u -1) \ from\ \ 1 \ to\ 0=1/2 \) nice work...

OpenStudy (turingtest):

\[\frac12\int\ln udu=\lim_{r\to\infty}\left.\frac12u(\ln u-1)\right]_0^r\]thanks, means a lot from you :D

OpenStudy (turingtest):

oh yeah after the sub it's from 1 to 0, right...

OpenStudy (anonymous):

\(u(\ln u-1)\) or \(u \ln u-1\)

OpenStudy (anonymous):

second is right ha?

OpenStudy (anonymous):

im wondering \[\lim_{u \rightarrow 0} u \ln u =0\]

OpenStudy (anonymous):

Integration by parts \(u=p^2\) and \(dv=pe^{-p^2}\) so \(v=\frac{1}{2} e^{-p^2}\) and we have \[I=\frac{p^2}{2} e^{-p^2}]_0^\infty-\int\limits_{0}^{\infty} p e^{-p^2} dp=1/2\]

OpenStudy (turingtest):

so it works by parts after directly all, nice :) I've done integrals very much like this before, I should've seen that at least my solution was novel though :D

OpenStudy (anonymous):

:)